OpenStudy (anonymous):

Mrs Perry shares out 15 biscuits between Gemma and Zak in the ratio 1:2 Howmany biscuits does each child get Gemma gets...... Biscuits and Zak get........ Biscuits

OpenStudy (anonymous):

Gemma gets 5 biscuits and Zak gets 10

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

1x + 2x = 15 3x = 15 3x/3 = 15/3 x = 5 Gemma = 1x, Zak = 2x therefore Gemma = 5 and Zak = 10

OpenStudy (anonymous):

mr castle shARES OUT 40 marbles between emily and asif in the ratio 3:1 . how many does each child get?

OpenStudy (anonymous):

okay you solve this one

OpenStudy (anonymous):

Emily = ? and Asif = ?

OpenStudy (anonymous):

3x+1x=4x 4x=40 x=40/4=10 x=10 emily= 3x10=30 asif=1x10=10

OpenStudy (anonymous):

coorect

OpenStudy (anonymous):

thanku

OpenStudy (anonymous):

welc